OpenStudy (nptweezy):

Solve For X –x + 10 – 8x = 17 –2x

OpenStudy (uri):

-x-8x+2x=17-10 -7x=7 x=7/-7 x=-1

OpenStudy (nptweezy):

@uri thanks

OpenStudy (uri):

yw :)